OpenStudy (anonymous):

Thomas wants to save money for a vacation. Thomas invests $1,500 in an account that pays an interest rate of 6.75%. How many years will it take for the account to reach $7,300? Round your answer to the nearest hundredth

OpenStudy (anonymous):

5800 = 1500 x 6.75 x t / 100

OpenStudy (amistre64):

1500(1.0675)^t=7300

OpenStudy (amistre64):

t = log(73/15) ---------- log(1.0675)

OpenStudy (amistre64):

log(73 / 15) / log(1.0675) = 24.2256916

OpenStudy (amistre64):

24.22 or maybe .23

OpenStudy (anonymous):

can you show me how you get that?

OpenStudy (amistre64):

sure..... me or him? lol

OpenStudy (anonymous):

bt i used simple interest

OpenStudy (anonymous):

u used compound interest amistre

OpenStudy (anonymous):

amistre

OpenStudy (amistre64):

i did :)

OpenStudy (amistre64):

the principle investment is 1500; and you mulitply that by( 1+.0675)^t

OpenStudy (anonymous):

but which approach is correct?

OpenStudy (amistre64):

1500(1.0675)^t = 7300 ; /1500 both sides (1.0675)^t = 7300/1500 = 73/15

OpenStudy (amistre64):

take the log of base (1.0675) to each side to get: log1.0675(1.0675^t) = log1.0675(73/15) / factor the ^t t log1.0675(1.0675) = log1.0675(73/15) logB(B) = 1 so....... t = log1.0675(73/15)

OpenStudy (amistre64):

change of base allows us to modify the right side to something we can use on a calculator

OpenStudy (amistre64):

logB(x) = log(x)/log(B)

OpenStudy (amistre64):

t = log(73/15)/log(1.0675)

OpenStudy (anonymous):

thank you for clearing it up for me!

OpenStudy (amistre64):

:) no prob, i figure in 3 days ill forget it anyways so its good to write it down lol

OpenStudy (anonymous):

lol yeah that's true. i just posted another question also. just need to know if i got the right answer & used the right equation for once!